# Forests of Uncertaint(r)ees: Using tree-based ensembles to estimate probability distributions of future conflict

### v1

@@ Abstract
Predictions of fatalities from violent conflict on the PRIO-GRID-month (pgm) level are characterized by high levels of uncertainty, limiting their usefulness in practical applications. We discuss the two main sources of uncertainty for this prediction task, the nature of violent conflict and data limitations, embedding this in the wider literature on uncertainty quantification in machine learning. We develop a strategy to quantify uncertainty in conflict forecasting, shifting from traditional point predictions to full predictive distributions. Our approach compares and combines multiple tree-based classifiers and distributional regressors in a custom auto-ML setup, estimating distributions for each pgm individually. We also test the integration of regional models in spatial ensembles as a potential avenue to reduce uncertainty. The models are able to consistently outperform a suite of benchmarks derived from conflict history in predictions up to one year in advance, with performance driven by regions where conflict was observed. With our evaluation, we emphasize the need to understand how a metric behaves for a given prediction problem, in our case characterized by extremely high zero-inflatedness. While not resulting in better predictions, the integration of smaller models does not decrease performance for this prediction task, opening avenues to integrate data sources with less spatial coverage in the future.
@@ 1 Introduction
With the advent of fine-grained conflict data and the increasing availability of computational resources over the last decade and a half, significant efforts have been made to design conflict forecasting systems to warn policy-makers of emerging dangers and enable preventative action. These systems generally provide estimates of the probability of conflict occurrence or point estimates of conflict-related fatalities (see Rød et al. (2024) for a comparison of existing systems). However, Hegre et al. (2025) highlight two key problems limiting the utility of both such point predictions: First, they only reflect the most likely scenario. Second, they do not provide any confidence estimates for the predicted value. Importantly, both issues are closely linked, as the latter becomes more important the less likely the underlying scenario is. The proposed solution is to estimate a predictive distribution, which allows for the quantification of uncertainties.
@@
The sources of uncertainty in conflict prediction can be broken down into uncertainties related to the nature of conflict, and uncertainties related to the (lack of) data used in conflict prediction. The former includes its complexity and related challenges of translating explanation to prediction (Ward et al., 2010), its relative rarity (Mueller and Rauh, 2022), its changing nature (Bowlsby et al., 2020; Hegre et al., 2021), and resulting doubts regarding the systematicity and therefore predictability of conflict (Chadefaux, 2017). The latter includes various data-related challenges in the field, such as concerns on quality, availability, and low resolution of data on potential predictors (e.g. Cederman and Weidmann, 2017; Mueller and Rauh, 2018), and a number of biases in conflict event datasets, many related to its reliance on news media (Weidmann, 2015; Althaus et al., 2022; Bazzi et al., 2022; Öberg and Yilmaz, 2025; Raleigh et al., 2023). Integrating this into a much larger body of research on uncertainty quantification (UQ) in machine learning (ML), one can also view these issues in terms of aleatoric and epistemic uncertainty, with the former resulting from inherently stochastic processes and therefore irreducible, and the latter related to data selection and the modeling process (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). While it can be debated how much of the apparently stochastic nature of conflict (Chadefaux, 2017) is in fact the product of data issues, from a statistical perspective the various methods to quantify uncertainty can be loosely summarized as estimating the predictive distribution for a given observation (Gruber et al., 2025). This aligns with the approach proposed by Hegre et al. (2025) for conflict forecasts.
@@
We design and evaluate a ML approach including UQ on the PRIO-GRID-month level (pgm, Tollefsen et al., 2012), combining various tree-based models. We specifically incorporate two key sources of data-related uncertainty: First, we rely on algorithms which are able to estimate distributions natively and for each pgm individually, accounting for local differences in data generation processes and conflict mechanisms. Second, we employ probabilistic hurdle ensembles to account for systematic biases between the recording of events and the recording of fatalities (Hegre et al., 2022b; Lacina, 2006). We select combinations of either Random Forest (Breiman, 2001) or XGBoost (Chen and Guestrin, 2016) classifiers with Distributional Random Forest (Cevid et al., 2022), Quantile Regression Forest (Meinshausen, 2006), or NGBoost (Duan et al., 2019) regression models through a custom AutoML [^1] AutoML refers to the automation of some or all components of machine learning, such as model or hyperparameter selection. For more information, see e.g. Hutter et al. (2019). setup. Additionally, we explore an avenue to address data availability issues and regionally varying conflict mechanisms by testing the integration of multiple regional models with limited geographical scope in a spatial ensemble. As a result, data coverage requirements for individual models are reduced, thus allowing for the incorporation of region-specific datasets. This yields three model specifications: a “global”-only model covering the whole area of interest, a “local”-only model combining all regional models, and a global-local model based on the best-performing individual components of each.
@@
Our evaluation shows that all three approaches generally outperform benchmarks across three distributional metrics with only very few exceptions. The global model and the global-local combined model score very similarly, while the local model performs only slightly worse. However, in absolute terms the differences in scores are marginal, prompting further investigation. Since the zero-inflatedness of our target greatly reduces the informative range of our scores, we design an experiment with simulated data to examine the impact of varying the accuracy of distributional predictions on the Continuous Ranked Probability Score (CRPS), our main metric. Our results show that the marginal differences in absolute scores observed are likely the consequence of sizable differences in prediction quality.
We complement the standard aggregate evaluation approach with a ranking-based approach. We group the predictions by matching pgms with countries, evaluate each country group individually, rank the model and benchmarks according to their performance, and compare the mean ranks across all countries for each year and metric. While this approach disadvantages our model compared to the benchmarks for countries without violence, since our distributional approach assigns a non-zero base probability for violence in any context, it outperforms the benchmarks in countries where violence occurs, which is arguably more relevant to users of early-warning systems.
@@ 2 Sources of Uncertainty in Conflict Prediction
Upon closer inspection, predicting violence with reasonable certainty is a daunting task. The many challenges can broadly be grouped into two categories: the nature of conflict and its determinants, and the characteristics, biases and availability of the data used. The uncertainty resulting from these challenges means that point predictions, i.e. predictions without uncertainty estimates, often struggle to beat even simple heuristics on common evaluation metrics (Vesco et al., 2022).
@@
Decades of research into the causes of armed conflict has identified a number of contributing factors from opportunity and feasibility to motivation and grievances (e.g. Fearon and Laitin, 2003; Collier et al., 2008; Collier, 2004; Blattman and Miguel, 2010), with differences between individuals and social groups (Humphreys and Weinstein, 2008; Østby, 2008; Cederman et al., 2013). However, this does not automatically mean these factors can be used to make reliable predictions for the future (Ward et al., 2010; Chadefaux and Schincariol, 2025). While they are known to increase the risk of conflict, no combination of these factors can be considered a sufficient condition for conflict, with very similar contexts resulting in widely different outcomes in practice. The complex nature of conflict together with the unpredictability of actors involved has even led some to question to what extent conflict is even systematic enough to be predicted (Chadefaux, 2017). The challenge is exacerbated by the relative rarity of conflict, especially at the pgm-level [^2] See 3.1: From 1990 to 2017, UCDP records violence in less than 0.4% of all pgms., which means in practice there are very few examples of conflict that can be used to derive very complex patterns from (Mueller and Rauh, 2022; Hegre et al., 2025). Uncertainty is further increased by continuously evolving situations on the ground and larger geopolitical shifts, which can affect underlying risk patterns and limit the validity of historic insights or change the conditions that might lead to violence (Cederman and Weidmann, 2017; Chadefaux, 2017; Bowlsby et al., 2020; Hegre et al., 2021).
@@
Second, the data used to capture both conflict and the various potential explanatory factors suffers from multiple shortcomings and biases. For many socio-economic and political risk factors of conflict, data varies widely in quality (Cederman and Weidmann, 2017; Chadefaux, 2017; Murphy et al., 2024), e.g. due to difficulties and expenses connected to measurement, or deliberate misrepresentation by governments. For many countries, subnational and subyearly data is also either not available or lags too far behind reality to be useful for forecasting tasks at the pgm-level. Additionally, data on structural risk factors often shows too little variance for meaningful predictions regarding the timing of future conflict (Mueller and Rauh, 2018; Chadefaux and Schincariol, 2025), while other data is not available at all or only in limited contexts. Consequently, the only fine-grained, high-variance features included often are conflict history features, which tend to have an outsized impact in terms of predictive power (Hegre et al., 2021, 2022a; Mueller and Rauh, 2022; Chadefaux and Schincariol, 2025). On the one hand, this corresponds to the well-documented issue of conflict recurrence (Collier et al., 2003), on the other hand conflict likely also proxies many unobserved factors that have already resulted in violence. Existing forecasts therefore do much better at discovering locations at risk than predicting when conflict is most likely to erupt (Mueller and Rauh, 2018; Bazzi et al., 2022).
@@
However, the conflict event datasets that forecasts rely on for the prediction target and the most important feature set also suffer from several biases and shortcomings themselves. Mostly based on news reporting, both the accuracy of the information contained and the selection of events covered are potential sources of bias (Earl et al., 2004; Althaus et al., 2022). While not perfect, information on the events ultimately recorded has been evaluated as reasonably accurate, with errors regarding exact location and number of fatalities within acceptable boundaries for the pgm prediction problem (Weidmann, 2015). In contrast, it is often hard to determine what portion of violence is reported at all and to what extent it is biased exactly, given the difficulty of establishing an accurate ground truth in a conflict context (Price and Ball, 2015; Althaus et al., 2022). Selection biases generally depend on the widely varying judgments of newsworthiness and (partisan) preferences by a particular source (Davenport and Ball, 2002; Baum and Zhukov, 2015; Dietrich and Eck, 2020), while also impacted by logistical factors such as the possibility for information to reach reporters (Weidmann, 2016; Croicu and Kreutz, 2017). In sum, this likely results in considerable variance in data quality across different contexts.
@@ 2.1 Uncertainty Quantification
While UQ has received only limited attention in conflict prediction, a large body of research discusses UQ in machine learning settings. Uncertainty can be conceptually split into aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty is caused by inherently stochastic processes and therefore irreducible, while epistemic uncertainty is the result of uncertainties related to the modeling process and can theoretically be minimized by additional data and better (fitted) models - although this may not always be possible in practice (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). Methods to quantify this uncertainty can be loosely summarized as methods to estimate the predictive distribution, depending on the problem resulting either in probabilistic classification or prediction intervals (Cheng et al., 2023; Gawlikowski et al., 2023; Tyralis and Papacharalampous, 2024; Gruber et al., 2025), with many existing approaches rooted in Bayesian analysis (Berger, 1985; Lampinen and Vehtari, 2001).
@@
Distinguishing between the types of uncertainty is often part of UQ approaches (Abdar et al., 2021; Hüllermeier and Waegeman, 2021; Gruber et al., 2025). However, Gruber et al. (2025) argue that data-related uncertainties blur the boundaries between aleatoric and epistemic uncertainty, with missing or biased data - a key challenge for conflict prediction - resulting in increased aleatoric uncertainty in practice, despite this being a theoretically solvable problem. In addition to what is discussed above, it could further be argued that the seemingly stochastic nature of conflict (Chadefaux, 2017) is simply the result of missing data, e.g. on the motivations of key actors, rather than an inherently stochastic process. Hence, instead of trying to disentangle these, focussing on the predictive distribution for UQ as proposed by the VIEWS team (Hegre et al., 2025) can be viewed as a sufficient first step in line with common practices for UQ in ML (Tyralis and Papacharalampous, 2024).
@@
With these considerations in mind, we design an ML-based approach to conflict prediction incorporating UQ. In principle, ML is well-suited to capture the complexity of conflict through its ability to account for non-linear and interaction effects between multiple variables (Cederman and Weidmann, 2017). We use tree-based models, which often outperform the more popular deep learning approaches on tabular data (Grinsztajn et al., 2022). While deep learning approaches are in principle more capable in big data applications, their potential has so far not materialized in the field of conflict forecasts, likely due to the relative rarity of conflict occurrence. Our UQ approach specifically addresses two main data-related sources of uncertainty:
@@
First, we exclusively use models with the ability to natively output uncertainty estimates. Importantly, given the potential of similar situations leading to widely different outcomes locally and over time, we focus on algorithms with the ability to predict distributions for each pgm individually rather than estimating uncertainty globally.
@@
Second, we implement a hurdle approach to distinguish between zero and non-zero observations. This assumes differences in distributions on either side of a threshold, which is arguably the case in conflict data due to a combination of biases in the data-generation process as well as differences between the determinants of the intensity of conflict and the determinants of the occurrence of conflict (Lacina, 2006; Hegre et al., 2022b). The two-stage approach focusing first on the likely location of violence before predicting its intensity is also well suited to capture the spatial clustering and diffusion patterns of armed conflict (Buhaug and Gleditsch, 2008; Schutte and Weidmann, 2011; Racek et al., 2025).
@@
Additionally, to work around some of the data availability challenges, we explore the selective combination of “local” and “global” models and its impact on model performance. While local predictions have been found to suffer from some of the same issues as the field of conflict forecasting as a whole (Bazzi et al., 2022), their integration into larger spatial ensembles is yet untested. Our contribution thus paves the way for the selective inclusion of locally available data sources into larger forecasting systems.

### v2

@@ Abstract
Predictions of fatalities from violent conflict on the PRIO-GRID-month (pgm) level are characterized by high levels of uncertainty, limiting their usefulness in practical applications. We discuss the two main sources of uncertainty for this prediction task, the nature of violent conflict and data limitations, embedding conflict prediction in the wider literature on uncertainty quantification in machine learning. Based on this, we develop a strategy to quantify uncertainty in conflict forecasting, shifting from traditional point predictions to full predictive distributions. Our approach combines multiple tree-based classifiers and distributional regressors in a custom AutoML setup, estimating distributions for each pgm individually. We also test the integration of regional models in spatial ensembles as a potential avenue to reduce uncertainty by lowering data requirements and accounting for systematic differences between conflict contexts. The models are able to consistently outperform a suite of benchmarks derived from conflict history in predictions up to one year in advance. Marginal differences in model-wide metrics emphasize the need to understand their behavior for a given prediction problem, in this case characterized by extremely high zero-inflatedness. Adressing this, we compliment our evaluation with a simulation experiment, which demonstrates that our models reflect meaningful performance improvements, which can be traced back to conflict-affected regions. Lastly, we show that the integration of regional models does not decrease performance, opening avenues to integrate additional data sources in the future.
@@ 1 Introduction
With the advent of fine-grained conflict data and the increasing availability of computational resources over the last decade and a half, significant efforts have been made to design conflict forecasting systems to warn policy-makers of emerging dangers and enable preventative action. These systems generally provide probability estimates for conflict occurrence or point estimates of conflict-related fatalities (see Rød et al. (2024) for a comparison of existing systems). However, Hegre et al. (2025) highlight two key problems limiting the utility of both types of point predictions: First, they only reflect the most likely outcome. Second, they do not provide any confidence estimates for the predicted value. Importantly, both issues are closely linked, as the latter becomes more important the less likely the underlying scenario is. Their proposed solution is to estimate a predictive distribution, which allows for the quantification of uncertainties.
@@
The sources of uncertainty in conflict prediction can be broken down into uncertainties related to the nature of conflict, and uncertainties related to the (lack of) data used in conflict prediction. The former includes the complexity of conflict and related challenges of translating explanation to prediction (Ward et al., 2010), its relative rarity (Mueller and Rauh, 2022), its changing nature (Bowlsby et al., 2020; Hegre et al., 2021), and resulting doubts regarding the systematicity and therefore predictability of conflict (Chadefaux, 2017). The latter includes various data-related challenges in the field, such as concerns over quality, availability, and low resolution of data on potential predictors (e.g. Cederman and Weidmann, 2017; Mueller and Rauh, 2018), and a number of biases in conflict event datasets, many related to its reliance on news media (Weidmann, 2015; Althaus et al., 2022; Bazzi et al., 2022; Öberg and Yilmaz, 2025; Raleigh et al., 2023). Integrating this into a much larger body of research on uncertainty quantification (UQ) in machine learning (ML), these issues can also be viewed in terms of aleatoric and epistemic uncertainty, with the former resulting from inherently stochastic processes and therefore irreducible, and the latter related to data selection and the modeling process (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). While there is room for debate about how much of the apparently stochastic nature of conflict (Chadefaux, 2017) is in fact the product of data issues, from a statistical perspective the various methods to quantify uncertainty can be loosely summarized as estimating the predictive distribution for a given observation (Gruber et al., 2025). This aligns with the approach proposed by Hegre et al. (2025) for conflict forecasts.
@@
We contribute to the understanding of uncertainty in conflict prediction by designing and evaluating a ML approach including UQ on the PRIO-GRID-month level (pgm, Tollefsen et al., 2012), combining various tree-based models. We specifically incorporate two key sources of data-related uncertainty: First, we rely on algorithms which are able to estimate distributions natively and for each pgm individually, accounting for local differences in data generation processes and conflict mechanisms. Second, we employ probabilistic hurdle ensembles to account for systematic differences between the recording of events and the recording of fatalities (Hegre et al., 2022b; Lacina, 2006). We select combinations of either Random Forest (Breiman, 2001) or XGBoost (Chen and Guestrin, 2016) classifiers with Distributional Random Forest (Cevid et al., 2022), Quantile Regression Forest (Meinshausen, 2006), or NGBoost (Duan et al., 2019) regression models through a custom AutoML [^1] AutoML refers to the automation of some or all components of machine learning, such as model or hyperparameter selection. For more information, see e.g. Hutter et al. (2019). setup. Additionally, we explore an avenue to address data availability issues and regionally varying conflict mechanisms by integrating multiple regional models with limited geographical scope in a spatial ensemble. If successeful, data coverage requirements for individual models are reduced to the corresponding region, thus allowing for the incorporation of region-specific datasets. This yields three model specifications: a “global”-only model covering the whole area of interest, a “local”only model combining all regional models, and a global-local model based on the best-performing individual components of each.
@@
Our evaluation shows that all three approaches outperform benchmarks across three distributional metrics with only very few exceptions. The global model and the global-local model score very similarly, while the local model performs only slightly worse. However, in absolute terms the differences in scores are marginal, prompting further investigation. Since the zero-inflatedness of our target greatly reduces the informative range of our scores, we design an experiment with simulated data to examine the impact of varying accuracy and noise in the predictions on our three metrics. The results show that the small differences in absolute scores observed likely reflect sizable differences in prediction quality for this problem. We complement the model-wide evaluation with a ranking-based approach, comparing the mean ranks based on performance across all countries. While this approach disadvantages our model compared to the benchmarks for countries without violence, when we restrict the evaluation to country-year instances with any observed violence – arguably the cases in which predictive performance is most relevant – our models clearly outperform the benchmarks.
@@ 2 Sources of Uncertainty in Conflict Prediction
Upon closer inspection, predicting violence with reasonable certainty is a daunting task. The many challenges can broadly be grouped into two categories: the nature of conflict and its determinants, and the characteristics, biases, and availability of the data used. The uncertainty resulting from these challenges means that point predictions, i.e. predictions without uncertainty estimates, often struggle to beat even simple heuristics on common evaluation metrics (Vesco et al., 2022).
@@
Decades of research into the causes of armed conflict has identified a number of contributing factors from opportunity and feasibility to motivation and grievances (e.g. Fearon and Laitin, 2003; Collier et al., 2008; Collier, 2004; Blattman and Miguel, 2010), with differences between individuals and social groups (Humphreys and Weinstein, 2008; Østby, 2008; Cederman et al., 2013). However, this does not automatically mean these findings translate into reliable predictions for the future (Ward et al., 2010; Chadefaux and Schincariol, 2025). While they are known to increase the risk of conflict, no combination of these factors can be considered a sufficient condition for conflict, with very similar contexts resulting in widely different outcomes in practice. The complex nature of conflict together with the unpredictability of actors involved has even led some to question to what extent conflict is even systematic enough to be predicted (Chadefaux, 2017). The challenge is exacerbated by the relative rarity of conflict, especially at the pgm-level, [^2] See 3.1: From 1990 to 2017, UCDP records violence in less than 0.4% of all pgms. which means in practice there are very few instances of conflict that can be used to derive very complex patterns from (Mueller and Rauh, 2022; Hegre et al., 2025). Uncertainty is further increased by continuously evolving situations on the ground and larger geopolitical shifts, which can affect underlying risk patterns and limit the validity of historic insights or change the conditions that might lead to violence (Cederman and Weidmann, 2017; Chadefaux, 2017; Bowlsby et al., 2020; Hegre et al., 2021).
@@
Exacerbating the problem, the data used to capture both conflict and the various potential explanatory factors suffers from multiple shortcomings and biases. For many socio-economic and political risk factors of conflict, data varies widely in quality (Cederman and Weidmann, 2017; Chadefaux, 2017; Murphy et al., 2024), e.g. due to difficulties and expenses connected to measurement, or deliberate misrepresentation by governments. For many countries, subnational and subyearly data is also either not available or lags too far behind reality to be useful for forecasting tasks at the pgm-level. Additionally, data on structural risk factors often shows too little variance for meaningful predictions regarding the timing of future conflict (Mueller and Rauh, 2018; Chadefaux and Schincariol, 2025), while other data does not exist at all or only in limited contexts. Consequently, the only fine-grained, high-variance features available often are conflict history features, which tend to have an outsized impact in terms of predictive power (Hegre et al., 2021, 2022a; Mueller and Rauh, 2022; Chadefaux and Schincariol, 2025). On the one hand, this corresponds to the well-documented issue of conflict recurrence (Collier et al., 2003), on the other hand conflict likely also proxies many unobserved factors that have already resulted in violence. Current forecasting systems therefore do much better at discovering locations at risk than predicting when conflict is most likely to erupt (Mueller and Rauh, 2018; Bazzi et al., 2022).
@@
However, the conflict event datasets that forecasts rely on, not only as the most important feature set but also for the prediction target, suffer from several biases and shortcomings themselves. Mostly based on news reporting, both the accuracy of the information contained and the selection of events covered are potentially affected (Earl et al., 2004; Althaus et al., 2022). While not perfect, information on the events ultimately recorded has been evaluated as reasonably accurate, with errors regarding exact location and number of fatalities within acceptable boundaries for the pgm prediction problem (Weidmann, 2015). In contrast, it is often hard to determine what portion of violence is reported at all and to what extent it is biased exactly, given the difficulty of establishing an accurate ground truth in a conflict context (Price and Ball, 2015; Althaus et al., 2022). Selection biases generally depend on the widely varying judgments of newsworthiness and (partisan) preferences by a particular source (Davenport and Ball, 2002; Baum and Zhukov, 2015; Dietrich and Eck, 2020), while also impacted by logistical factors such as the possibility for information to reach reporters (Weidmann, 2016; Croicu and Kreutz, 2017). In sum, this likely results in considerable variance in data quality across different contexts.
@@ 2.1 Uncertainty Quantification
While UQ has received only limited attention in conflict prediction, a large body of research discusses UQ in machine learning settings. Uncertainty can be conceptually split into aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty is caused by inherently stochastic processes and therefore irreducible, while epistemic uncertainty is the result of uncertainties related to the modeling process and can theoretically be minimized by additional data and better (fitted) models – although this may not always be possible in practice (Hüllermeier and Waegeman, 2021; Gruber et al., 2025). Methods to quantify this uncertainty can be loosely summarized as methods to estimate the predictive distribution, resulting either in probabilistic classification or prediction intervals (Cheng et al., 2023; Gawlikowski et al., 2023; Tyralis and Papacharalampous, 2024; Gruber et al., 2025), with many existing approaches rooted in Bayesian analysis (Berger, 1985; Lampinen and Vehtari, 2001).
@@
Distinguishing between the types of uncertainty is often part of UQ approaches (Abdar et al., 2021; Hüllermeier and Waegeman, 2021; Gruber et al., 2025). However, Gruber et al. (2025) argue that data-related uncertainties blur the boundaries between aleatoric and epistemic uncertainty, with missing or biased data – a key challenge for conflict prediction – resulting in increased aleatoric uncertainty in practice, despite this being a theoretically solvable problem. Given the data-related challenges discussed above, at least some seemingly stochastic aspects of conflict (Chadefaux, 2017) may simply be the result of missing data, e.g. on the motivations of key actors, rather than an inherently stochastic process. Hence, focusing on the predictive distribution for UQ as proposed by the VIEWS (Violence & Impacts Early-Warning System) team (Hegre et al., 2025) without trying to disentangle these types of uncertainty constitutes a sufficient first step in line with common practices for UQ in ML (Tyralis and Papacharalampous, 2024).
@@
With these considerations in mind, we design a ML-based approach to conflict prediction incorporating UQ. In principle, ML is well-suited to capture the complexity of conflict through its ability to account for non-linear effects and interaction effects between multiple variables (Cederman and Weidmann, 2017). We use tree-based models, which often outperform the more popular deep learning approaches on tabular data (Grinsztajn et al., 2022). While deep learning approaches are in principle more capable in big data applications, their potential has so far not materialized in the field of conflict forecasts, likely connected to the relative rarity of conflict occurrence. Our UQ approach specifically addresses two main data-related sources of uncertainty:
@@
First, we exclusively use ML algorithms with the ability to natively output uncertainty estimates. Importantly, given the potential of similar situations leading to widely different outcomes locally and over time, we focus on algorithms with the ability to predict distributions for each pgm individually rather than estimating uncertainty globally.
@@
Second, we implement a hurdle approach to distinguish between zero and non-zero observations. This assumes differences in distributions on either side of a threshold, which is arguably the case in conflict data due to a combination of biases in the data-generation process as well as differences between the determinants of the occurrence of conflict and the determinants of the intensity of conflict (Lacina, 2006; Hegre et al., 2022b). The two-stage approach focusing first on the likely location of violence before predicting its intensity is also well suited to capture the spatial clustering and diffusion patterns of armed conflict (Buhaug and Gleditsch, 2008; Schutte and Weidmann, 2011; Racek et al., 2025).
@@
Additionally, to work around some of the data availability challenges, we explore the selective combination of “local” and “global” models and its impact on model performance. While local predictions have been found to suffer from some of the same issues as of conflict forecasting as a whole (Bazzi et al., 2022), their integration into larger spatial ensembles is yet untested. Our contribution thus paves the way for the selective inclusion of locally available data sources into larger forecasting systems.